in how many ways can a research team of 5 members be formed from a group consisting of 3 chemist and 7 physicist if
a. there is no restriction in the selection?
b. the team must include exactly 2 chemist?
c. the team must include at most 3 physicists?

Answer by Boreal(15235)

a. This is 10 C 5 =252 ways
b. 3C2*7C3, since there have to be 3 physicists. This is 3*35=105
c. At most 3 physicists includes the first of 105
2 physicists and 3 chemists, which is 3C3*7C2=1*21=21 ways
1 physicist and 4 chemists doesn't work, because there are only 3 chemists.
therefore, 126 ways. ANSWER